Solve the equation by completing the square.
9y^2 -12y -2 = 0

Answer by stanbon(75887)

Solve the equation by completing the square.
9y^2 -12y -2 = 0
---
9y^2-12y = 2
---
9(y^2-(4/3)y+?) = 2+9(?)
---
9(y^2-(4/3)y+(2/3)^2) = 2+9(2/3)^2
---
9(y-(2/3))^2 = 2+4
(y-(2/3))^2 = 2/3
------------
y-(2/3) = +/- sqrt(2/3)
---
y = (2/3)+sqrt(2/3) or y = (2/3)-sqrt(2/3)
==============================================
Cheers,
Stan H.
========